OpenStudy (anonymous):

If a light has a energy of 342.8J/mol of photons, find it's energy in J of just 1 photon

OpenStudy (anonymous):

n = N/L so n = 1/ 6,022*10^23 mol-1 and then just divide given number of 342,8 with calculated number